7. The curve y = (x + 2)2 has a minimum point at?

8. The coordinates of the lowest point on the curve: y = x2 - 4x + 1 are?


1
Expert's answer
2021-02-24T07:42:52-0500

7. If branches of a parabola a directed upwards and it has roots "a" and "b", then the parabola has minimum point at "x =\frac{a+b}{2}". "y=(x+2)^2" is a parabola with repeated root "-2", which means it has minimum at "-2".

8. For a parabola "y=ax^2+bx+c" we can use equation "x_m=-\frac{b}{2a}" for x coordinate of minimum point. Parabola "y = x^2 - 4x + 1" has minimum point at "x_m=-\frac{-4}{2*1}=2" and it's y coordinate is "y_m=2^2-4*2+1=-3" . So, the minimum point is "(2,-3)"